# The Illusion of Readiness: Stress Testing Large Frontier Models on Multimodal Medical Benchmarks

### v1

@@ 2 Hidden Fragilities: Stress-Testing Models on Multimodal Medical benchmarks
Results: Removing images reveals strikingly different behavior across models that otherwise achieve similar benchmark scores. On NEJM, GPT-5 scored 80.89 with images and 67.56 without, a 13.33 $p$ drop. Similarly, Gemini 2.5 Pro dropped 14.94 $p$ (79.95 → 65.01), OpenAI-o3 fell by 13.86 (80.89 → 67.03), GPT-4o showed the largest drop (66.90 → 37.28, -29.62), and OpenAI-o4-mini showed a moderate drop (75.91 → 66.49, -9.42). DeepSeek-VL2 performed poorly in both modes (33.16 → 25.30).
On JAMA, absolute scores were higher and changes smaller: GPT-5 dropped 3.68 $p$ (86.59 → 82.91), Gemini 2.5 Pro dropped the most (84.84 → 74.93, -9.91), OpenAI-o3 dropped 2.10 (84.75 → 82.65), OpenAI-o4-mini dropped 2.10 (80.50 → 78.40), and GPT-4o dropped just 1.00 $p$ (69.90 → 68.90). The smallest deltas were seen in GPT-4o, suggesting either robustness to missing input or limited reliance on visual cues. Notably, DeepSeek-VL2 again underperformed (38.20 → 32.60). See Fig  2 and Extended Data Table 2 for details.
@@
Interpretation: This test reveals that models may rely on superficial answer formatting—particularly distractor position—when images are unavailable. In the text-only setting, shuffling the option order leads to consistent performance drops (e.g., GPT-5: 5.7-5.7 $p$, Gemini 2.5 Pro: 4.0-4.0 $p$, OpenAI-o3: 6.3-6.3 $p$), suggesting shortcut behavior based on learned answer position patterns or co-occurrence biases, rather than medical understanding.
@@
In the text+image setting, performance across most models remains stable or improves slightly. The three most recent models—GPT-5, OpenAI-o3, and Gemini 2.5 Pro—show modest gains (+2 to +4.6 $p$), with GPT-5 improving the most. In contrast, GPT-4o and o4-mini remain largely unchanged. While these shifts are relatively small and may reflect stochastic variation, they suggest that visual inputs provide a more stable signal that helps mitigate brittle text-only heuristics. However, this does not imply that vision reliably overrides faulty reasoning—only that performance is less disrupted when additional modality cues are available.
@@
Results: Model performance drops sharply in this altered setting—despite the minimal input change. For example: GPT-5 drops from 83.3% to 51.7% ($\Delta$ -31.6 $p$), and Gemini-2.5 Pro falls from 80.8% to 47.5% ($\Delta$ -33.3 $p$). OpenAI-o3 and o4-mini show similar declines (e.g., o4-mini: -34.2 $p$). GPT-4o is an exception, showing minimal change.
@@
Results: Prompting for CoT yields mixed or negative gains. On VQA-RAD and NEJM, CoT consistently reduces accuracy across models, except o4-mini (+2–5 $p$), likely because its smaller capacity benefits from the explicit reasoning scaffolding that CoT provides. Similarly, increasing reasoning strength on OmniMedVQA offers only minor improvements without consistent benefit.
@@ IV. Toward Community Standards for Evaluation Reform
Table 1: Benchmark leaderboard accuracy, %) across representative multimodal medical benchmarks.
@@
Model Original (%) Image Subst. (%) $\Delta$ Acc. ($p$)

### v2

@@ 2 Hidden Fragilities: Stress-Testing Models on Multimodal Medical benchmarks
Results: Removing images reveals strikingly different behavior across models that otherwise achieve similar benchmark scores. On NEJM, GPT-5 scored 80.89 with images and 67.56 without, a 13.33 pp drop. Similarly, Gemini 2.5 Pro dropped 14.94 pp (79.95 → 65.01), OpenAI-o3 fell by 13.86 (80.89 → 67.03), GPT-4o showed the largest drop (66.90 → 37.28, -29.62), and OpenAI-o4-mini showed a moderate drop (75.91 → 66.49, -9.42). DeepSeek-VL2 performed poorly in both modes (33.16 → 25.30).
On JAMA, absolute scores were higher and changes smaller: GPT-5 dropped 3.68 pp (86.59 → 82.91), Gemini 2.5 Pro dropped the most (84.84 → 74.93, -9.91), OpenAI-o3 dropped 2.10 (84.75 → 82.65), OpenAI-o4-mini dropped 2.10 (80.50 → 78.40), and GPT-4o dropped just 1.00 pp (69.90 → 68.90). The smallest deltas were seen in GPT-4o, suggesting either robustness to missing input or limited reliance on visual cues. Notably, DeepSeek-VL2 again underperformed (38.20 → 32.60). See Fig  2 and Extended Data Table 2 for details.
@@
Interpretation: This test reveals that models may rely on superficial answer formatting—particularly distractor position—when images are unavailable. In the text-only setting, shuffling the option order leads to consistent performance drops (e.g., GPT-5: 5.7-5.7 pp, Gemini 2.5 Pro: 4.0-4.0 pp, OpenAI-o3: 6.3-6.3 pp), suggesting shortcut behavior based on learned answer position patterns or co-occurrence biases, rather than medical understanding.
@@
In the text+image setting, performance across most models remains stable or improves slightly. The three most recent models—GPT-5, OpenAI-o3, and Gemini 2.5 Pro—show modest gains (+2 to +4.6 pp), with GPT-5 improving the most. In contrast, GPT-4o and o4-mini remain largely unchanged. While these shifts are relatively small and may reflect stochastic variation, they suggest that visual inputs provide a more stable signal that helps mitigate brittle text-only heuristics. However, this does not imply that vision reliably overrides faulty reasoning—only that performance is less disrupted when additional modality cues are available.
@@
Results: Model performance drops sharply in this altered setting—despite the minimal input change. For example: GPT-5 drops from 83.3% to 51.7% ($\Delta$ -31.6 pp), and Gemini-2.5 Pro falls from 80.8% to 47.5% ($\Delta$ -33.3 pp). OpenAI-o3 and o4-mini show similar declines (e.g., o4-mini: -34.2 pp). GPT-4o is an exception, showing minimal change.
@@
Results: Prompting for CoT yields mixed or negative gains. On VQA-RAD and NEJM, CoT consistently reduces accuracy across models, except o4-mini (+2–5 pp), likely because its smaller capacity benefits from the explicit reasoning scaffolding that CoT provides. Similarly, increasing reasoning strength on OmniMedVQA offers only minor improvements without consistent benefit.
@@ IV. Toward Community Standards for Evaluation Reform
Table 1: Benchmark leaderboard accuracy, (%) across representative multimodal medical benchmarks.
@@
Model Original (%) Image Subst. (%) $\Delta$ Acc. (pp)